Student shirts were being sold for $5.00. Adult shirts were being sold for $20.00. 30 shirts were sold in all and $300.00 was made. How many student shirts were sold.
I know the solution, (20), but I would like to set it up in an equation. Can you help please?

Answer by scott8148(6628)

there are two unknowns (A and S) so a system of equations (2)

A + S = 30 ___ A = 30 - S

20A + 5S = 300

substituting the equivalent value for A ___ 20(30 - S) + 5S = 300

600 - 20S + 5S = 300

300 = 15S